OpenStudy (anonymous):

f(x) = e^(2x) + e^(−2x) Find an upper bound for the error in using the second degree Maclaurin polynomial of f to approximate f(0.5).

OpenStudy (anonymous):

I guess I have to use the Lagrange remainder formula, right? \[E _{n}(x) = f(x) - T _{n}(x) =\frac{ f ^{(n+1)}(c) }{ (n+1)! }(x-a)^{(n+1)}\] Where c lies between x and a

OpenStudy (anonymous):

yes just plug the rest in

OpenStudy (anonymous):

So here a=0 right?

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

Is my 3rd derivative correct? I got... \[8e ^{2x}-8e ^{-2x}\]

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

Cool, so then I get \[E _{2}(0.5)=\frac{ e ^{2c} -e ^{-2c}}{ 6 }\]

OpenStudy (anonymous):

yes! im surprised u asked for help cuz u get it

OpenStudy (anonymous):

Lol I don't know what to do next :D

OpenStudy (anonymous):

that is f(x)

OpenStudy (anonymous):

i see what u do now

OpenStudy (anonymous):

u have to

OpenStudy (anonymous):

replace x with .5

OpenStudy (anonymous):

What if I was trying to find the lower bound?